OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -negative one fourthx - 5

OpenStudy (anonymous):

please help !!!!!!!!!!!!!!!!!!! desparate !!!!!!!!!!!!

OpenStudy (anonymous):

@adventuretimefreak7 @camrontheking @cordell10 @jjuden

OpenStudy (anonymous):

@YanaSidlinskiy @arabpride @skullpatrol

OpenStudy (anonymous):

please help sombody

OpenStudy (arabpride):

any answer choices?

OpenStudy (anonymous):

yea hold on

OpenStudy (anonymous):

y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (anonymous):

here

OpenStudy (arabpride):

y = -1/4x which one of the options are negative reciprocals to the equation that'll help u figure out ur answer *^.^*

OpenStudy (anonymous):

sorry im terrible at this what do you mean

OpenStudy (arabpride):

well to find the perpendicular line, u have to find the negative reciprocal of -1/4x

OpenStudy (anonymous):

so its would be A

OpenStudy (arabpride):

well okay.. here is how it works: negative reciprocal: finding the reciprocal which would be 4/-1x then u wud multiply it by -1 that's why its called 'negative' reciprocal, nd u wud get -4/1x which = -4x that wud mean, ur equation would have to have a -4x for ur slope so.... either A or B

OpenStudy (anonymous):

ok which do you think a or b

OpenStudy (arabpride):

i thnk its y = -4 + 3